OpenStudy (anonymous):

5√6x+8√14x-9√6x+18√14x

OpenStudy (jiteshmeghwal9):

\[5\sqrt{6x}-9\sqrt{6x}+8\sqrt{14x}+18\sqrt{14x}\]solve all these like terms :)

OpenStudy (anonymous):

thanks

OpenStudy (jiteshmeghwal9):

yw :)